# ELDeR: Getting Efficient LLMs through Data-Driven Regularized Layer-wise Pruning

### v1

@@ Abstract
The deployment of Large language models (LLMs) in many fields is largely hindered by their high computational and memory costs. Recent studies suggest that LLMs exhibit sparsity, which can be used for pruning. Previous pruning methods typically follow a prune-then-finetune paradigm. Since the pruned parts still contain valuable information, statically removing them without updating the remaining parameters often results in irreversible performance degradation, requiring costly recovery fine-tuning (RFT) to maintain performance. To address this, we propose a novel paradigm: first apply regularization, then prune. Based on this paradigm, we propose ELDeR: Getting Efficient LLMs through Data-Driven Regularized Layer-wise Pruning. We multiply the output of each transformer layer by an initial weight, then we iteratively learn the weights of each transformer layer by using a small amount of data in a simple way. After that, we apply regularization to the difference between the output and input of the layers with smaller weights, forcing the information to be transferred to the remaining layers. Compared with direct pruning, ELDeR reduces the information loss caused by direct parameter removal, thus better preserving the model’s language modeling ability. Experimental results show that ELDeR achieves superior performance compared with powerful layer-wise structured pruning methods, while greatly reducing RFT computational costs. Since ELDeR is a layer-wise pruning method, its end-to-end acceleration effect is obvious, making it a promising technique for efficient LLMs.
@@ 1 Introduction
Large language models (LLMs) have made remarkable progress in various tasks and fields, showcasing their strong capabilities [1, 2, 3, 4]. As models grow larger, the increasing number of parameters results in substantial computational and memory demands, posing major challenges to the practical deployment of LLMs. There is an urgent need for techniques that can compress model size without compromising performance.
The main methods for reducing model size include pruning [5, 6, 7], quantization [8, 9], distillation [10], and low-rank decomposition [11]. This study primarily concentrates on pruning, an effective and widely adaptable technique that is orthogonal to other model compression methods. Pruning methods are typically divided into two main types: unstructured pruning [12, 13] and structured pruning [14, 15]. In contrast to unstructured pruning, structured pruning enables recovery fine-tuning (RFT) for specific downstream tasks without the need for specialized hardware. It can also achieve significant end-to-end acceleration without the need for specially designed hardware [16].
Despite advancements, existing structured pruning methods have some limitations. They all follow the paradigm of first prune channels or layers based on a designed metric, and then performing RFT. However, they overlook that important information may also exist in the pruned parts [17, 18, 19], and directly removing them without updating the remaining parameters often results an irreversible decline in model performance. Therefore, obtaining satisfactory performance with pruned models typically requires a large amount of data for RFT [14], increasing computational overhead. Moreover, high sparsity can cause the model’s performance to collapse, limiting the acceleration effect.
@@
To address these limitations, we present a new pruning paradigm that first apply regularization and followed by pruning. The regularization process dynamically transfers crucial information from the layers to be pruned to the remaining layers, thus minimizing the information loss resulting from direct parameter removal. This paradigm can maintain good model performance without RFT. Based on this paradigm, we propose ELDeR, a simple and effective data-driven regularized layer-wise pruning method. The comparison of the ELDeR with existing layer-wise pruning approaches is shown in Figure 1. Specifically, the process of ELDeR is divided into the following steps: (1) select a small amount of data for learning layer weights and regularization, (2) iteratively learn the layer weights, (3) apply regularization to the difference between the outputs and inputs of the layers to be pruned, (4) prune. First, we randomly choose a small subset of data from commonly used benchmark datasets. Since the selected data is small, the computational costs in learning the layer weights and regularization process can be greatly minimized. Second, we iteratively learn the weights of each layer by adding their $\ell_{1}$-norm in the loss function. Third, we apply regularization ($\ell_{1}$-norm or $\ell_{2}$-norm) to the difference between the outputs and inputs of the layers with smaller weights, forcing important information to be transferred to the remaining layers which significantly reduces the performance decline caused by parameter removal. Thus the model can maintain good performance even at high pruning ratios, thereby providing significant end-to-end acceleration. Finally, we prune the layers with smaller weights. Extensive experiments demonstrate that ELDeR substantially outperforms current layer-wise pruning methods in terms of perplexity and accuracy across different pruning ratios, while also significantly improving end-to-end acceleration. The main contributions of ELDeR are summarized as follows:
@@
  • Propose A New Paradigm: ELDeR reduces information loss caused by direct parameter removal by sequentially learning layer weights, applying regularization ($\ell_{1}$-norm or $\ell_{2}$-norm), and performing pruning. This approach helps preserve model performance without requiring for RFT.
@@
  • High Performance: ELDeR outperforms previous layer-wise structured pruning methods in both generation and zero-shot tasks. The pruned model demonstrates a considerable acceleration compared to the dense model.
  • Low Overhead: By using only a small amount of data for learning layer weights and regularization, ELDeR significantly reduces the overhead of RFT.
@@
Figure 1: A comparison between existing layer-wise structured pruning methods and ELDeR. Deeper blue layer represents greater performance impact, the taller cylinder represents larger data volume.
@@ 2 Related Works
2.1 Pruning Methods
The process of removing unnecessary weights, known as pruning, has proven to be an effective method for reducing the complexity of deep neural networks over the past few decades [20, 21, 22, 23]. Pruning methods can generally be divided into two categories: unstructured pruning [24, 25, 26] and structured pruning [5, 27, 28].
Unstructured pruning methods eliminate individual weights according to their significance. For instance, Magnitude [22, 29] assumes larger weights are more important, Wanda [13] uses the product of the layer weights and activations as the criterion, and SparseGPT [12] uses the weight squared divided by the inverse Hessian matrix. Their main advantage is flexibility. However, their drawbacks include the need for specialized hardware to accelerate [30], the requirement for extreme sparsity to achieve considerable acceleration [31], and are not suitable for RFT on downstream tasks.
Structured pruning methods, including channel-wise pruning [15] and layer-wise pruning [32]. Channel-wise pruning methods typically create a metric to assess the importance of channels in the parameter matrixs of LLMs, and then remove the less important ones. Layer-wise pruning methods treat entire layers as the basic units for pruning. For instance, SLEB [16], iteratively remove entire transformer layers by evaluating their impact on the model’s final loss, ShortGPT [33] thinks high similarity between layers means redundancy, LaCo [27] uses layer collapse to prune, Shortened LLaMA [34] prune layers in one-shot based on their importance. However, these methods have a major limitation: even less important channels or layers may still contain valuable information, and directly pruning them often results in significant performance degradation. To overcome this, we explore the possibility of "shifting" important information from the pruned parts to the remaining parts, which could greatly maintain model performance.
@@ 2.2 Regularization
Regularization is commonly used in machine learning [35, 36, 37], including preventing model overfitting [38] and feature selection [39]. The $\ell_{1}$-norm typically promotes sparsity by driving certain coefficients to zero, while the $\ell_{2}$-norm encourages smoother solutions [40]. Both can significantly affect the distribution pattern of the data [22, 41]. Inspired by this, we propose a new pruning paradigm, in which the regularization process could transfer important information from the pruned parameter space to the remaining parts of the model, thus improving performance.
@@ 3 Methodology
In this section, we provide a detailed introduction to ELDeR. The overall ELDeR algorithm is summarized in Algorithm 1. ELDeR mainly consists of four steps:
  • Select Data: Select a small amount of data to learn the layer weights and apply regularization.
  • Learn Layer Weights: Iteratively learn the weights of each layer by adding their $\ell_{1}$-norm to the loss function.
  • Regularization: Apply regularization to the difference between the output and input of layers with smaller weights, facilitating information transfer.
  • Pruning: Removes the layers that were regularized in the previous step.
3.1 Select Data
The data is selected from widely used datasets, including Alpaca [42], WikiText-2 [43], PTB [44], and C4 [45]. For instance, approximately 128 samples are randomly selected from the WikiText-2 training dataset for learning layer weights and applying regularization.
3.2 Learn Layer Weights
For clarity, we first define the relevant notations. $\mathbf{W}$ represents the initial model. Let $p$ be the pruning ratio, indicating that $p\%$ of model parameters will be pruned. The number of layers in the initial model is denoted by $l$. The model hidden size is $d$. Let $\mathbf{X}\in\mathbb{R}^{b\times n\times d}$ represents the input data embedding, where $b$ is the batch size and $n$ is the number of tokens. The input to the $i$th layer is denoted as $\mathbf{X_{\text{in}}^{i}}\in\mathbb{R}^{b\times n\times d}$, and the output from the $i$th layer is denoted as $\mathbf{X_{\text{out}}^{i}}\in\mathbb{R}^{b\times n\times d}$. A learnable weight $S[i]$ is assigned to the $i$th transformer layer, and the set of all layer weights is denoted as $S$.
Algorithm 1 ELDeR algorithm.
1:Input: selected data: $\mathbf{X}$, number of layers in LLM: $l$, initial model $\mathbf{W}$, number of layers to prune: $n$, set of layer weights: $S$, set of layers to be pruned: $P$, norm type: flag.
@@ 3.2 Learn Layer Weights
According to Algorithm 1, we initialize the weight of each layer to 1. As shown in Figure 2, the output of each layer is scaled by its associated weight before being passed as input to the subsequent layer. To learn the weight of each layer, we employ the input data embedding $\mathbf{X}$. When pruning $n$ layers, $\mathbf{X}$ is repeatedly used as input in each iteration. The objective function defined in Equation 1 comprises two parts: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$, and the sum of the $\ell_{1}$-norm of all layer weights, $\lambda_{1}$ balances the two components, $l_{1}$ is the number of layers in the current model that are not masked.
@@
We formalize the minimization of the objective function in Equation 1 as an optimization problem. Since $\ell_{1}$-norm is not differentiable, backpropagation (BP) can’t be used directly, we need to transform the problem using Proposition 3.1.
Proposition 3.1.
(If the loss function includes the $\ell_{1}$-norm, it can also be solved using BP. Proof in Appendix A). The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.
$\min\quad||x||_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (2)
@@
The optimization problem in Equation 1 can be equivalently transformed into a constrained formulation in Equation 3. After transformation, the objective function is differentiable, and can be solved by BP.
@@
3.3 Regularization
After completing learning the layer weighs, we further lighten the impact of the pruned layers on the
@@ 3.3 Regularization
final model output by applying regularization to them. The approach is simple and straightforward: we use the input data embedding $\mathbf{X}$ as input and apply a one-shot regularization on the difference between the output and input of each layer in the pruning set $P$. For example, in Figure 2, the $i$th layer in $P$ and should be regularized, we just add the $\ell_{1}$-norm or $\ell_{2}$-norm of $\mathbf{X_{out}^{i}-X_{in}^{i}}$ to the loss function. This encourages the information to be redistributed from the pruned layers to the remaining ones, thereby reducing the amount of information retained in the pruned layers and significantly minimizing the performance degradation caused by subsequent pruning.
The objective function defined in Equation 4 comprises two components: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$ and the regularization loss, $\lambda_{2}$ balances the two components. When the $\ell_{1}$-norm is used, the equivalent constrained optimization problem transformed using Proposition 3.1 is shown in Equation 5. This formulation can be directly solved using BP.
@@
$\min_{W,Y_{i}}$ $\quad\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{2}\sum_{i=0}^{|P|-1}\mathbf{1}^{T}\mathbf{Y_{i}}\mathbf{1}^{T}$ (5)
@@ 3.4 Pruning
After applying the regularization, we directly remove the transformer layers in the set $P$, which is a straightforward process.
@@ 4 Experiments
This section introduces experimental setup (4.1) and analyzes the effectiveness of ELDeR from the following aspects: performance comparison (4.2), acceleration 4.3, robustness under different pruning ratios (4.4), dependency on different datasets (4.5), low overhead (4.6), and ablation study (4.7), choice of learning layer weights (4.8), impact of regularization (4.9).
@@ 4.1 Experimental Setup
Implementation: All methods are implemented in PyTorch [46], using the Hugging Face Transformers library [47]. All experiments are conducted on 80GB NVIDIA A100 GPUs. For fairness, we use llm-eval-harness [48] to evaluate the pruned models. More details are in Appendix B.
Datasets: We evaluated on generation and zero-shot tasks. For generation task, following prior work[15], we evaluate the model’s perplexity on WikiText-2 [43] test set. For zero-shot task, we evaluate on PIQA [49], WinoGrande [50], HellaSwag [51], ARC-e and ARC-c [52]. In Section 4.5 we use a randomly selected subset of data from Alpaca [42], WikiText-2 [43], PTB [44], and C4 [45].
Models: The models pruned using ELDeR include the LLaMA models (LLaMA2-7B, LLaMA2-13B, LLaMA3-8B) [1, 53], OPT models (OPT-2.7B, OPT-13B) [54], and Phi-2 [55].
Baselines: We evaluate ELDeR against competitive layer-wise structured pruning methods: SLEB [16], ShortGPT [33], LaCo [27], Shortened LLaMA (PPL version) [34]. We also compared with powerful unstructured pruning methods like Magnitude [22], Wanda [13], and SparseGPT [12] in Appendix D.
Evaluation Metrics: The performance on generation task is measured by perplexity, a well-established and robust metric [56], while zero-shot tasks performance is evaluated using accuracy [57]. The acceleration effects are represented by throughput and latency [16].

### v2

# Two-Stage Regularization-Based Structured Pruning for LLMs

@@ Abstract
The deployment of large language models (LLMs) is largely hindered by their large number of parameters. Structural pruning has emerged as a promising solution. Prior structured pruning methods directly remove unimportant parameters based on certain metrics, which often causes knowledge loss and necessitates extensive retraining. To overcome this, we introduce a novel pruning method TRSP: Two-Stage Regularization-Based Structured Pruning for LLMs. Specifically, we multiply the output of each transformer layer by an initial learnable weight and iteratively learn these weights by adding their $\ell_{1}$-norm as a regularization term to the loss function, serving as the first-stage regularization. Subsequently, we apply additional regularization to the difference between the output and input of layers with smaller weights, encouraging the shift of knowledge to the preserved layers. This serves as the second-stage regularization. TRSP retains more knowledge and better preserves model performance than direct parameter elimination. Through extensive experimentation we show that TRSP outperforms strong layer-wise structured pruning methods without requiring retraining. As a layer-wise pruning method, it delivers notable end-to-end acceleration, making it a promising solution for efficient LLM deployment.
Two-Stage Regularization-Based Structured Pruning for LLMs
Mingkuan Feng1∗, Jinyang Wu1∗ Siyuan Liu2thanks: Equal Contribution., Shuai Zhang1, Ruihan Jin1, Feihu Che1, Pengpeng Shao1, Zhengqi Wen1, Jianhua Tao1 1Tsinghua University, 2Peking University
@@ 1 Introduction
Large language models (LLMs) have made remarkable progress in natural language processing Yang et al. (2024a); Wu et al. (2024); Guo et al. (2025). However, their large scale makes real-world deployment challenging. There is an urgent need for techniques that can enhance the compactness and computational efficiency of LLMs while preserving their language modeling capabilities.
Structured pruning is a method used to simplify neural networks by removing unnecessary or redundant parameters Xia et al. (2024); An et al. (2024); Feng et al. (2025). Structured pruning is categorized into channel-wise pruning Ma et al. (2023); Ashkboos et al. (2024) and layer-wise pruning Song et al. (2024). Channel-wise pruning operates at the row or column level of parameter matrices. Layer-wise pruning operates at the level of entire transformer layers thereby offering a more simple approach compared to channel-wise pruning Chen et al. (2024); Men et al. (2024).
However, existing layer-wise pruning methods have a certain limitation. They consistently first compute the importance of each transformer layer using a designed criteria, prune unimportant layers, and then fine-tune the pruned model to compensate for performance degradation caused by pruning. However, even unimportant layers can hold valuable knowledge Dettmers et al. (2022); Yin et al. (2024); An et al. (2025). This sequential process of selecting and then directly pruning layers does not handle the important knowledge contained in the layers that are to be pruned, resulting in its direct loss. The performance drop requires substantial retraining for recovery, leading to considerable computational overhead Ma et al. (2023).
@@
To address this, we present TRSP that first apply two-stage regularization and then prune. The first regularization process iteratively learns layer weights. The second regularization process dynamically transfers valuable knowledge from the layers to be pruned to the remaining layers in advance, greatly reducing the knowledge loss caused by pruning. The comparison of TRSP with existing layer-wise pruning approaches is shown in Figure 1. First, we sample a small portion of data from standard benchmark datasets randomly. Given the limited scale of the selected data, the computational overhead incurred during the two-stage regularization is significantly reduced. Second, we multiply the output of each transformer layer by an initial learnable weight and iteratively learn these weights by incorporating their $\ell_{1}$-norm as a regularization term in the loss function. This serves as the first stage regularization. Third, we apply regularization ($\ell_{1}$-norm or $\ell_{2}$-norm) to the difference between the outputs and inputs of the layers with smaller weights, forcing important knowledge to be transferred to the remaining layers which significantly reduces the performance decline caused by parameter removal. Thus the model can maintain good language modeling capability. This serves as the second stage regularization. Finally, we prune the layers with smaller weights. Comprehensive experiments demonstrate that TRSP substantially outperforms strong layer-wise pruning methods in generation tasks and zero-shot tasks across different pruning ratios, while also significantly improving end-to-end acceleration. The main contributions of TRSP are summarized as follows:
@@
  • Retention of Knowledge: TRSP reduces knowledge loss by progressively applying two-stage regularization and performing pruning. This approach helps preserve model performance without requiring for retraining.
@@
  • Effectiveness: TRSP outperforms strong layer-wise pruning methods in generation and zero-shot tasks. The pruned model demonstrates a considerable acceleration.
  • Minimal Cost: The data required for two-stage regularization is minimal and TRSP is retraining free after pruning.
@@
Figure 1: A comparison between existing layer-wise structured pruning methods and TRSP . Deeper blue layer represents greater performance impact, the taller cylinder represents larger data volume.
@@ 2 Related Works
2.1 Model Pruning
Model Pruning aims to improve model efficiency by sparsification or parameter removal LeCun et al. (1989); Hassibi et al. (1993); Han et al. (2015); Liu et al. (2017). Several studies employ unstructured Kurtic et al. (2022); Zhang et al. (2024); Xu et al. (2024) and structured pruning Xia et al. (2024); Yang et al. (2024b); Gao et al. (2024b).
Unstructured pruning zeros individual neurons according to their importance such as SparseGPT Frantar and Alistarh (2023), SpQR Dettmers et al. (2024), Pruner-Zero Dong et al. (2024), and Wanda Sun et al. (2024) . Their main advantage is flexibility. However, they need dedicated hardware to accelerate Xia et al. (2023), and are not able to retrain on downstream tasks.
Structured pruning methods can be categorized by granularity into channel-wise pruning Ashkboos et al. (2024) and layer-wise pruning Kim et al. (2024a). Channel-wise pruning methods create a metric to assess the significance of channels in the parameter matrices of LLMs, and then remove the less significant ones. Layer-wise pruning methods treat entire layers as the basic units for pruning. For instance, SLEB Song et al. (2024) iteratively remove entire transformer layers by evaluating their impact on the model’s final loss, ShortGPT Men et al. (2024) thinks high similarity between layers means redundancy, LaCo Yang et al. (2024b) uses layer collapse to prune, Shortened LLaMA Kim et al. (2024b) prune layers in one-shot based on their importance. In this paper, we focus on layer-wise pruning. Prior layer-wise pruning approaches suffer from the drawback that the less important layers may still carry critical knowledge, and pruning them often causes knowledge loss. Discovering a way to reshape knowledge distribution prior to pruning could potentially alleviate knowledge loss.
@@ 2.2 Regularization
In machine learning, regularization plays a vital role in controlling overfitting Santos and Papa (2022) and identifying informative features Tibshirani (1996), and has been extensively studied Hoerl and Kennard (1970); Poggio et al. (1987); Balestriero et al. (2022). The $\ell_{1}$-norm tends to enforce compact representations by eliminating certain parameters, whereas the $\ell_{2}$-norm favors stability and continuity Boyd and Vandenberghe (2004). Both can both alter the underlying structure and representation of the data Han et al. (2015); Tao et al. (2023). Inspired by this insight, regularization can be leveraged to migrate critical knowledge from pruned layers to preserved layers, thereby enhancing model performance.
@@ 3 Methodology
The complete TRSP procedure is outlined in Algorithm 1 and Figure 1. The TRSP framework involves four key stages: (1) Prepare data: Select a small amount of data for the following two-stage regularization. (2) Learn layer weights: Iteratively learn the weights of each layer by incorporating their $\ell_{1}$-norm as a regularization term in the loss function. (3) The second stage regularization: Apply regularization to the difference between the output and input of layers with smaller weights, facilitating knowledge transfer. (4) Pruning: Removes the layers that were regularized in the previous step.
Algorithm 1 TRSP algorithm.
1:Input: selected data: $\mathbf{X}$, number of layers: $l$, initial model $\mathbf{W}$, number of layers to prune: $n$, layer weights: $S$, set of pruned layers: $P$, norm type: flag.
@@
3.1 Prepare Data
The data is drawn from standard benchmarks, including Alpaca Taori et al. (2023), WikiText-2 Merity et al. (2016), PTB Marcus et al. (1993), and C4 Raffel et al. (2020). For example, 128 instances are randomly drawn from the WikiText-2 training set to for layer weight learning via regularization and the second stage regularization.
3.2 Learn Layer Weights
This is the first stage regularization. To better understand our paper, we first define some notations. $\mathbf{W}$ represents the initial model. Let $p$ be the pruning ratio, indicating that $p\%$ of model layers will be pruned. The number of layers in the initial model is $l$. The model hidden size is $d$. Let $\mathbf{X}\in\mathbb{R}^{b\times n\times d}$ represents the data embedding, where $b$ is the batch size and $n$ is the number of tokens. The input to the $i$th layer is denoted as $\mathbf{X_{\text{in}}^{i}}\in\mathbb{R}^{b\times n\times d}$, and the output from the $i$th layer is denoted as $\mathbf{X_{\text{out}}^{i}}\in\mathbb{R}^{b\times n\times d}$. A learnable weight $S[i]$ is assigned to the $i$th transformer layer, and the set of all layer weights is denoted as $S$.
According to Algorithm 1, we initialize the weight of each layer to 1. As shown in Figure 2, the output of each layer is scaled by its associated weight before being passed as input to the next layer. To learn the weight of each layer, we employ the input data embedding $\mathbf{X}$. When pruning $n$ layers, $\mathbf{X}$ is repeatedly used as input in each iteration. The objective function in Equation 1 comprises two parts: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$, and the sum of the $\ell_{1}$-norm of all layer weights, $\lambda_{1}$ balances the two components, $l_{1}$ is the number of layers in the current model that are not masked.
@@ 3.2 Learn Layer Weights
The process of minimizing the function in Equation 1 is treated as an optimization task. Since $\ell_{1}$-norm is not differentiable, backpropagation (BP) can’t be used directly, we need to transform the problem using Proposition 3.1.
Proposition 3.1
(If the objective function contains an $\ell_{1}$ regularization term, it can still be optimized using BP. Proof in Appendix A). The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.
$\min\quad||x||_{1}\hskip 5.0pt\iff\hskip 5.0pt\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (2)
@@
The objective function in Equation 1 can be reformulated as an equivalent constrained problem in Equation 3. After transformation, the objective function is differentiable, and can be solved by BP.
@@
3.3 The Second Stage Regularization
@@
After learning the layer weighs, we further lighten the impact of the pruned layers on the final model output by applying regularization to them. The approach is simple and straightforward: we use the data embedding $\mathbf{X}$ as input and apply a one-shot regularization on the difference between the output and input of each layer in the pruning set $P$. For example, in Figure 2, the $i$th layer in $P$ and should be regularized, we just add the $\ell_{1}$-norm or $\ell_{2}$-norm of $\mathbf{X_{out}^{i}-X_{in}^{i}}$ to the loss function. This encourages the knowledge to be redistributed from the pruned layers to the remaining ones, thereby reducing the amount of knowledge retained in the pruned layers and significantly minimizing the performance degradation caused by subsequent pruning.
The objective function in Equation 4 comprises two parts: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$ and the regularization loss, $\lambda_{2}$ balances the two components. When the $\ell_{1}$-norm is used, the equivalent constrained optimization problem transformed using Proposition 3.1 is shown in Equation 5. This formulation can be directly solved using BP.
@@
$\min_{W,Y_{i}}$ $\quad\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{2}\sum_{i=0}^{|P|-1}\mathbf{1}^{T}\mathbf{Y_{i}}\mathbf{1}$ (5)
@@ 3.4 Pruning
After applying the regularization, we directly remove the transformer layers in the set $P$.
@@ 4 Experiments
This section introduces experimental setup (4.1) and analyzes the effectiveness of TRSP from the following aspects: performance comparison (4.2), acceleration 4.3, robustness under different pruning ratios (4.4), dependency on different datasets (4.5), low overhead (4.6), and ablation study (3), choice of learning layer weights (4.8), impact of regularization (4.9).
@@ 4.1 Experimental Setup
Datasets: We evaluated on generation and zero-shot tasks. For generation task, following prior workAshkboos et al. (2024), we evaluate the model’s perplexity on WikiText-2 Merity et al. (2016) test set. For zero-shot task, we evaluate on PIQA Bisk et al. (2020), WinoGrande Sakaguchi et al. (2021), HellaSwag Zellers et al. (2019), ARC-e and ARC-c Clark et al. (2018). In Section 4.5 we randomly selected data from Alpaca Taori et al. (2023), WikiText-2 Merity et al. (2016), PTB Marcus et al. (1993), and C4 Raffel et al. (2020).
Implementation: All methods are developed in PyTorch Paszke et al. (2019), leveraging the Hugging Face Transformers library Wolf (2019). Experimental evaluations are carried out on 80GB NVIDIA A100 GPUs. Pruned models are evaluated with llm-eval-harness Gao et al. (2024a). More details are in Appendix B.
Evaluation Metrics: The ability on generation task is evaluated by perplexity, a well-established and robust metric Yao et al. (2022). Zero-shot tasks ability is evaluated using accuracy Dong et al. (2024). Acceleration is measured by throughput and latency Song et al. (2024).
Models: The models include the Phi-2 Javaheripi et al. (2023), OPT models (OPT-2.7B, OPT-13B) Zhang et al. (2022), and LLaMA models (LLaMA2-7B, LLaMA2-13B, LLaMA3-8B) Touvron et al. (2023); Grattafiori et al. (2024).
Baselines: We compare TRSP with strong layer-wise structured pruning methods: SLEB Song et al. (2024), ShortGPT Men et al. (2024), LaCo Yang et al. (2024b), Shortened LLaMA (PPL version) Kim et al. (2024b).